OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates (3, 5pi/3)

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8

jimthompson5910 (jim_thompson5910):

hint: x = r*cos(theta) y = r*sin(theta)

OpenStudy (anonymous):

is x 3/2?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so y must be -3/2? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

nope

OpenStudy (anonymous):

oh wait no it is -3sqrt 3/2???

jimthompson5910 (jim_thompson5910):

much better \[\Large y = \frac{-3*\sqrt{3}}{2}\]

OpenStudy (anonymous):

thanks!